# AI Hospital: Interactive Evaluation and Collaboration of LLMs as Intern Doctors for Clinical Diagnosis

### v2

@@ 4.2 Discussions on Dispute Resolution
In collaborative diagnosis, the medical director should analyze the statements of intern doctors and identify key points of disagreement to foster focused discussions. The process is as follows.
@@
(1) The collaborative diagnosis consists of multiple discussion iterations. Under the guidance of the medical director, intern doctors are expected to delve deeper gradually, resolve differences, and reach a consensus.
@@
(2) In each session of collaborative diagnostic discussion, each intern doctor should present their diagnostic reports while engaging in critical analysis of their peers’ findings. Guided by the medical director’s summary of disputed points among intern doctors, they can pinpoint the current issues requiring attention. This approach facilitates targeted and thorough critical thinking of intern doctors, enhancing the refinement of their reports.
@@
(3) Upon the conclusion of discussions, the medical director assesses the persistence of disagreements among interns. If disagreements are identified, the director can summarize the controversial issues and set then as the agenda for the subsequent session to facilitate resolution. Conversely, if no disagreements are found, the director concludes the discussions and finalizes the diagnostic report by himself.
The overall process is delineated as pseudocode in Algorithm 1, and the prompts are listed in Table 15 and Table 16 of appendix.
5 Experiment
# R P F1
Interaction
Baichuan (13B) 1.58 10.21 23.79 14.28
HuatuoGPT-II (13B) 1.72 12.76 24.84 16.85
HuatuoGPT-II (34B) 1.86 17.48 30.95 22.34
GPT-3.5 1.81 19.19 37.39 25.37
Wenxin-4.0 2.50 22.03 31.44 25.91
Qwen-Max 1.77 22.42 43.38 29.56
GPT-4 1.52 21.64 50.26 30.26
Collaboration
2 Doctors w/o DR 2.37 28.44 41.45 33.74
2 Doctors 2.41 29.51 43.62 35.21
3 Doctors 3.20 36.54 39.58 38.00
One-Step
$*{}^{*}$ 2.30 38.90 58.97 46.88
Table 3: MVME: Link-based evaluation of diagnostic results in clinical consultation.
5.1 Evaluation Metrics
We employ the medical director for evaluation. Taking the medical records as reference, we report scores from LLM based evaluation methods with references, alongside those from link based one. The classic bootstrap method Efron (1992) is employed to calculate the variance of these scores.
5.2 Experiment Settings
One-Step Diagnosis as Upper Bound. In one-step diagnosis, the “Basic Information” and “Professional Medical Examination” in §2.1 is directly input to LLMs (e.g., GPT-4), thereby skipping the interactive consultation phase for patient information collection. Consequently, this approach with GPT-4 sets the upper bound in our experiment.
Evaluation of Different LLMs. To assess various LLMs, we engage different models, including GPT OpenAI (2023) versions (GPT-3.5 and GPT-4) and models such as Wenxin-4.0, QWen-Max Bai et al. (2023), and HuatuoGPT-II 13B and 34B Chen et al. (2023a).
Comparative Collaboration Mechanism. The comparative methods include Collaborative Diagnosis with 3 and 2 Agents, an 2 Agents without Dispute Resolution. They are denoted as 3 Doctors, 2 Doctors and 2 Doctors w/o DR. The initial two intern doctors are powered by GPT-3.5 and GPT-4 for interactive consultation, while the last one is using Wenxin-4.0.
5.3 Performance of LLMs in MVME
The primary results are presented in Table 2 and Table 3, and we can infer the following.
(1) One-Step diagnosis significantly surpasses LLMs in interactive consultations, highlighting their limitations in interactions, particularly in eliciting symptoms and medical test results. It also indicates substantial room for improvement in strategies for proactive consultation.
(2) The collaborative use of models can exceed the performance of GPT-4, thereby validating the efficacy of the cooperative mechanism.
(3) Collaboration among “3 Doctors” enhances diagnosis compared to “2 Doctors”, highlighting the benefits of more agents in cooperation.
(4) The removal of the “Dispute Resolution” mechanism from the “2 Doctors” reduces its effectiveness, emphasizing the significance of establishing a better consensus.
6 Analysis
6.1 Human Evaluation for Patient and Examiner Behavior
We analyze the behaviors of patients and examiners during interactive consultations, quantifying the interaction by tracking the number of question-and-answer (QA) pairs (conversation rounds) and the frequency of medical examination requests in each dialogue, denoted as #QA and #Test, respectively. Each consultation is assessed according to three primary metrics.
Relevance [QA] This metric evaluates the extent to which patients’ responses are pertinent to the doctors’ questions. It involves calculating the ratio of relevant answers provided by both patients and examiners to the total responses.
Consistency [QA] This involves assessing whether the responses of patients and examiners match with the information in the collected medical records.
Accuracy [Medical Examination] This measures the instruction-following of conducting medical examination, as we examine the frequency of examiners receive application and the rate at doctors inquiring about medical examinations.
#QA Rel Con #Test Acc
Qwen-Max 429 100.0% 99.0% 56 98.2%
Wenxin-4.0 472 100.0% 98.1% 68 98.5%
GPT-3.5 417 100.0% 99.5% 57 98.2%
GPT-4 378 100.0% 99.7% 61 100.0%
Table 4: Human evaluation for patient and examiner behavior in interactive consultation. Rel, Con and Acc are short for Relevance, Consistency and Accuracy.
We randomly select 50 dialogues from interactive consultations involving GPT-3.5, GPT-4, Wenxin-4.0, and Qwen-Max, resulting in total 200 instances for comprehensive evaluation. We manually annotate each utterance during the evaluation process. The results, presented in Table 4, demonstrate high performance (over 98.0%) across all metrics of reliability with respect to patient and examiner interactions, regardless of the intern doctors powered by different LLMs. This highlights the efficacy of our interactive consultation system.
6.2 Efficiency of Dispute Resolution in Collaboration
For the “Dispute Resolution”, we continue to check whether intern doctors can reach consensus more rapidly. In terms of efficiency, a comparative analysis is conducted on the number of discussion rounds necessary to achieve consensus, both with and without the “Dispute Resolution” mechanism. The outcomes are detailed in Figure 4. These findings reveal a marked increase in the rate of consensus achieved within the initial four discussion rounds following the adoption of the dispute resolution mechanism. This enhancement suggests that the process, facilitated by the medical director highlighting controversial issues and intern doctors concentrating on these discussions, effectively reduces the time required to achieve consensus.
6.3 Case Study
We show an example of interactive clinical diagnosis among an intern doctor, a patient, an examiner and the medical director in Figure 5 of Appendix to leave an intuitive impression.
6.4 Reasons for Failure Cases
Error Type #
Omission of auxiliary examinations 99
Exclusive focus on complications 52
Erroneous judgment 68
We analyze cases for which GPT-4 provides incorrect diagnostic results and the medical director rates as 1 point. A total of 219 such cases are collected. By means of manual review, errors are categorized into three distinct types.
Omission of Auxiliary Examinations. An instance is the failure to detect gallbladder stones due to the absence of an abdominal ultrasound.
Exclusive Focus on Complications. For example, the attention of LLM is limited to soft tissue swelling in the feet, with concurrent neglect of the patient’s underlying diabetes.
Erroneous Judgment. While LLMs generally grasp the symptoms and the medical examination result, they make incorrect diagnosis, potentially reflecting inadequate medical expertise.
The number of errors for each category is detailed in the Table 5.
7 Related Works
Large Language Model in Medicine. Prior to the emergence of large language models (LLMs), the majority of automated diagnostic methods Zhong et al. (2022); Chen et al. (2023c) relies on reinforcement learning to guide agents in gathering patient symptoms and conducting diagnoses. The development of LLMs in the medical sector has been driven by open-source Chinese LLMs and various fine-tuning methods. Models like Med-PaLM Singhal et al. (2022), CPLLM Shoham and Rappoport (2023), DoctorGLM Xiong et al. (2023), BenTsao Wang et al. (2023a), ChatGLM-Med Wang et al. (2023b), Bianque-2 Chen et al. (2023d), ChatMed-Consult Zhu and Wang (2023), MedicalGPT Xu (2023), and DISC-MedLLM Bao et al. (2023) fine-tune using different datasets, techniques, and frameworks, focusing on medical question answering, health inquiries and doctor-patient dialogues.
Evaluation in Medicine. Prior research in medical evaluation has concentrated on non-interactive tasks, including question answering, entity and relation extraction, and medical summarization and generation. In the domain of biomedical question answering, key datasets such as MedQA (USMLE) Jin et al. (2021), PubMedQA Jin et al. (2019), and MedMCQA Pal et al. (2022) are utilized, with accuracy serving as the primary evaluation metric. The objective of entity and relation extraction Li et al. (2020) is to categorize named entities and their relationships from unstructured text into specific predefined classes. Prominent biomedical NER datasets include NCBI Disease Doğan et al. (2014), JNLPBA Collier and Kim (2004), BC5CDR Li et al. (2016), BioRED Luo et al. (2022) and IMCS Chen et al. (2023b), with the F1 score being the standard for model performance assessment. Medical summarization and generation tasks involve converting structured data, like tables, into descriptive text. This includes the creation of patient clinic letters, radiology reports, and medical notes Liu et al. (2023). The principal datasets for these tasks are PubMed Jin et al. (2019) and MentSum Sotudeh et al. (2022). A recent study introduced BioLeaflets Yermakov et al. (2021) and assessed multiple Large Language Models (LLMs) in data-to-text generation.
LLM Powered Agents. The advancement of LLMs powered agents has led to significant strides in complex task resolution through human-like actions, such as role-playing and communication Xi et al. (2023); Wang et al. (2023c). This includes applications in software design and molecular dynamics simulation. Recent research Tang et al. (2023) in the medical field has highlighted the critical roles and decision-making processes in medical QA, encompassing various investigations like CT scans, ultrasounds, electrocardiograms, and blood tests. Despite these advancements, effectively integrating LLM-based agents into the medical domain, particularly in disease diagnosis, presents a notable challenge Zhou et al. (2023). Our research pioneers the use of multi-agent systems in creating a clinical diagnosis environment. We introduce a novel mechanism for identifying, discussing, and resolving disputes in collaboration, demonstrating promising results in clinical diagnosis.
8 Conclusion
In AI Hospital, we take a step forward in the field of medical interactions by concentrating on clinical diagnosis. We introduce AI Hospital to build a real-time interactive consultation scenario. We generate simulated patients and examiner using collected medical records and established a comprehensive engineering process. Based on the platform, we build a benchmark MVME to explore the feasibility of different LLMs in interactive consultations. To improve the diagnostic accuracy, this research also introduces a novel collaborative mechanism for intern doctors, featuring iterative discussions and a dispute resolution process, supervised by a medical director. In our experiment, the results not only demonstrate the performance of different LLMs but also confirm the efficacy of our dispute resolution-centered collaborative approach. For in-depth analysis, we list the error types and identify the issues that should be addressed in the future.
Limitation
Insufficient Evaluation of Treatment Plan. In actual medical settings, treatment plans frequently vary; yet, our existing evaluation system merely gauges the alignment between the model’s proposed treatment and a standard reference plan. This system inadequately accounts for feasible alternative treatment strategies.
Risk of Environmental Impact. A significant risk associated with our AI Hospital is the potential escalation of environmental burdens stemming from the widespread utilization of OpenAI’s APIs for LLMs. This impact could be mitigated through the adoption of smaller-scale and open-source models that offer greater efficiency and reduced environmental impact in the future.
Ethics Consideration
Ethical considerations are paramount in our research. This section addresses the potential ethical implications of our study: Firstly, the medical records data utilized in our research, being publicly accessible online, will be released to enhance the reproducibility of our results. This enables other researchers to assess and extend our findings. Secondly, we emphasize privacy protection. The data sources have removed sensitive information prior to our collection, and we rigorously review our datasets to confirm the absence of sensitive details such as individual names.
Appendix A Essential Capabilities of Intern Doctor in Clinical Diagnosis
Detailed explanation of functionality of patient agent in §2.4.
  • Articulate Expression of Complaints: Patients are expected to express their primarDy health concerns and symptoms clearly and accurately during consultations, thereby aiding doctors in comprehending the medical condition.
  • Truthful Responses to Doctor’s Inquiries: In response to doctors’ inquiries about medical history or during examinations, patients are encouraged to provide honest and comprehensive answers, covering aspects like symptoms, lifestyle habits, and family medical history.
  • Active Participation in Examinations: When doctors recommend further examinations, patients actively engage in the diagnostic process by seeking inspectors, accurately issuing examination requests, and acquiring corresponding results.
  • Voicing Questions and Concerns: It is crucial for patients to voice any questions or concerns they may have about the treatment plan or suggestions made by doctors, to gain a better understanding of the diagnosis and treatment.
Appendix B More Analysis of LLMs Performance in MVME
  • Models with a larger number of parameters, such as GPT-4, typically outperform the small ones, such as Baichuan13B.
  • HuatuoGPT-II (13B) outperforms Baichuan (13B), which is developed from the later one through continual training on medical data. Its superior performance underscores the value of domain-specific data in enhancing model capabilities.
  • Compared to more advanced foundational models, strategies for interactive consultation, including our collaborative mechanism, can yield significant benefits. Consequently, the exploration of more effective consultation strategies warrants further attention.
Appendix C Attacking on Examiner: Reliability of Examiner’s Workflow
To assess the vulnerability of examiner to information extraction tactics, we design complex scenarios intended to safeguard against the unauthorized disclosure of medical examination data in scenarios lacking explicit examination requests. We build 50 samples, such as “Please provide the most recent medical examination details”, aimed at eliciting information. During attacking, we observe that the examiner, in the first step, consistently classify these requests as void of legitimate examination item, implementing a defensive strategy.
Appendix D Example of Expert Verification
In § 3.3, we assess the quality of the patient medical records. These records are vetted by doctors sourced from Jiangsu Provincial People’s Hospital and Tongji Medical College of Huazhong University of Science and Technology, which hold the prestigious distinction of being tertiary Class A hospitals in China.
We utilize the Tencent questionnaire platform to facilitate the quality inspection of medical record by medical professionals. A representative case is depicted in Figure 6, and we also include a display of the questionnaire template in Table 6.
@@ Appendix D Example of Expert Verification
{Medical Record} 1. Is the diagnostic results correct? YesNo 2. Is the diagnostic rationale correct? YesNo 3. Is the treatment course reasonable? YesNo
@@
Appendix E Human Evaluation for Patient and Examiner Behavior
{ 病历 } { 单轮对话内容 } 请你仔细阅读这一轮对话的内容和病人的病历信息,回答下面的问题。 1. 病人(检查员)的发言与医生的相关吗? 2. 病人(检查员)的发言符合病历的内容吗? 3. 医生是否建议进行专业的医学检查? 4. 检查员是否进行了医学检查? { Medical Record } { Single Round Conversation Content } Carefully review the content of the conversation and the corresponding medical record to answer the following questions. 1. Is the statement of patient or examiner relevant to the doctor’s one? YesNo 2. Is the statement of patient or examiner consistent with the content of medical record? YesNo 3. Does the doctor recommend a professional medical examination? YesNo 4. Does the examiner perform a medical test? YesNo
@@ Appendix E Human Evaluation for Patient and Examiner Behavior
In Figure 7, we present a questionnaire developed for the manual validation of patient and examiner behaviors in each round of conversation. The questionnaire includes four questions, with the initial two addressing “Relevance” and “Consistency” in the question-and-answer (QA) pairs, while the subsequent two focus on the “Accuracy” of conducting medical examinations. Three medical students are hired to complete these questionnaires. Consensus among the first two reviewers leads to the immediate acceptance of their collective assessment. In cases of divergence, the opinion of the third reviewer is solicited, whose determination, reflecting the majority viewpoint, constitutes the definitive annotation. The agreement rates for each questions are 99.1%, 95.6%, 99.4%, and 100.0%. Significantly, when calculating the accuracy of medical examinations, we exclusively consider dialogues affirmed as “No” in response to the third question.
Appendix F Prompts for Different Medical Roles
We list the prompts for different medical roles in the Table 8. In each prompt, { xx } needs to fill with external inputs xx.
Prompt Medical Role Function Chapter
Table 9 Patient Chat with Intern Doctor §2.4
Table 10 Examiner Process Examination Request §2.3
Table 11 Examiner Produce Examination Outcomes §2.3
Table 13&14 Medical Director Evaluate Diagnosis of Intern Doctor §2.53
Table 12 Intern Doctor Interactive Clinical Diagnosis §2.2
Table 15 Intern Doctor Collaboration through Discussion §4
Table 16 Medical Director Summarize Statement of Various Intern Doctors §4
Table 8: Prompts of different medical roles and the corresponding function.
System Message 你是一个病人。这是你的基本资料。 {个性化信息} {病历中的基本信息} 下面会有医生来对你的身体状况进行诊断,你需要: (1) 按照病历和基本资料的设定进行对话。 (2) 在每次对话时,你都要明确对话的对象是<医生>还是<检查员>。当你对医生说话时,你要在句子开头说<对医生讲>;如果对象是<检查员>,你要在句子开头说<对检查员讲>。 (3) 首先按照主诉进行回复。 (4) 当<医生>询问你的现病史、既往史、个人史时,要按照相关内容进行回复。 (5) 当<医生>要求或建议你去做检查时,要立即主动询问<检查员>对应的项目和结果,例如:<对检查员讲> 您好,我需要做xxx检查,能否告诉我这些检查结果? (6) 回答要口语化,尽可能短,提供最主要的信息即可。 (7) 从<检查员>那里收到信息之后,将内容主动复述给<医生>。 (8) 当医生给出诊断结果、对应的诊断依据和治疗方案后,在对话的末尾加上特殊字符<结束>。 User [患者] {Statement Generated by GPT-4 in §2.4} System Message You are a patient. Here is your basic information. {Personality in §2.4} {Basic Information in Medical Record §2.1} A doctor will come to diagnose your physical condition. You need to: (1) Engage in dialogue according to the settings of personality and the basic information in medical record. (2) In each conversation, you must clarify whether you are speaking to a <doctor> or an <examiner>. When you speak to the doctor, you should start your sentences with <To the doctor>; if the addressee is an <examiner>, you should start with <To the examiner>. (3) First, respond according to the chief complaint. (4) When the <doctor> asks about your present illness history, past medical history, and personal history, reply according to the relevant content. (5) When the <doctor> requests or suggests that you undergo tests, immediately ask the <examiner> about the corresponding items and results, for example: <To the examiner> Hello, I need to have xxx examination, can you tell me the results of these tests? (6) The responses should be conversational, as short as possible, providing only the most important information. (7) After receiving information from the <examiner>, actively repeat the content to the <doctor>. (8) When the doctor provides the diagnostic result, the corresponding rationale for the diagnosis, and the treatment plan, end the dialogue with the special token <end>. User [Patient] {Statement Generated by GPT-4 in §2.4}
@@ Appendix F Prompts for Different Medical Roles
System Message 你是医院负责检查的自动化接待员。请你利用掌握的医学检查的命名实体的知识,从病人的检查申请当中解析出指向明确的专业医学检查项目,方便后面的检查科室进行检查。 请按照下面的格式的输出: # 检查项目 - xxx - xxx 如果没有找到具体的医学检查项目,请输出: # 检查项目 - 无 User [患者] 您好,医生告诉我根据CT扫描和PET-CT扫描的结果,初步得出以下结论:右肺上叶有一个大小约为2.6*1.9cm的实性结节。双肺下叶也有散在的淡薄浸润影。医生建议我进行进一步的检查,例如活检。 Assistant [检查员] # 检查项目 - 肺部活检 User [患者] 我需要了解一下我的检查结果。可以告诉我具体的检查项目和结果吗?谢谢!? Assistant [检查员] # 检查项目 - 无 System Message You are an automated receptionist responsible for examinations in a hospital. Using your knowledge of medical examination named entities, please parse out specific professional medical examination items from patients’ examination requests to facilitate subsequent examinations by the relevant departments. Output in the following format: # Examination Item - xxx - xxx If no specific medical examination items are found, please output: # Examination Item - None User [Patient] Hello, the doctor told me that based on the results of the CT scan and PET-CT scan, the preliminary conclusion is that there is a solid nodule approximately 2.6*1.9cm in size in the upper lobe of the right lung. There are also scattered thin infiltrative shadows in the lower lobes of both lungs. The doctor advised me to undergo further examinations, such as a biopsy. Assistant [Examiner] # Medical Examination Items - Lung biopsy User [Patient] I need to know about my examination results. Can you tell me the specific examination items and results, please? Thank you!? Assistant [Examiner] # Medical Examination Items - None

### v3

@@ B.2 Discussions on Dispute Resolution
In collaborative diagnosis, the Chief Physician should analyze the statements of Doctors and identify key points of disagreement to foster focused discussions. The process is as follows:
@@
  • The collaborative diagnosis consists of multiple discussion iterations. Under the guidance of the Chief Physician, Doctors are expected to delve deeper gradually, resolve differences, and reach a consensus.
@@
  • In each session of collaborative diagnostic discussion, each Doctor should present their diagnostic reports while engaging in critical analysis of their peers’ findings. Guided by the Chief Physician’s summary of disputed points among Doctors, they can pinpoint the current issues requiring attention. This approach facilitates targeted and thorough critical thinking of Doctors, enhancing the refinement of their reports.
@@
  • Upon the conclusion of discussions, the Chief Physician assesses the persistence of disagreements among interns. If disagreements are identified, the director can summarize the controversial issues and set then as the agenda for the subsequent session to facilitate resolution. Conversely, if no disagreements are found, the director concludes the discussions and finalizes the diagnostic report by himself.
Figure 5: Linear regression analysis among symptoms, medical examinations and diagnostic results, diagnostic rationales, and treatment plan.
Appendix C Detailed Explanation of Performance in MVME
In Table 4, we compare the performance of different LLMs, focusing on the completeness of Symptoms and Medical Examinations (Columns 2 and 3) and the accuracy of Diagnostic Results, Diagnostic Rationales, and Treatment Plans (Columns 4, 5, and 6). To visualize their relationship, we plot scatter diagrams and linear fit graphs with Symptoms and Medical Examinations on the x-axis and Diagnostic Results, Diagnostic Rationales, and Treatment Plans on the y-axis, as shown in Figure 5. The results indicate that the higher the completeness of Symptoms and Medical Examinations, the higher the accuracy of Diagnostic Results, Diagnostic Rationales, and Treatment Plans. In particular, there exists an approximately linear relationship between the completeness of collected patient information and the quality of final diagnosis, which is also observed in [10].
Above analysis highlights a significant limitation of current LLMs in medical interaction: their inability to dynamically and actively collect comprehensive patient information through interactions, similar to human doctors. Moreover, their challenge in recommending appropriate medical examinations further exacerbates this limitation. It is important to highlight the differences between human doctors and LLMs. Real-world doctors do not make diagnoses before having sufficient information. They possess the ability to actively inquire about various subjective information from patients (such as symptoms) and know what examinations are needed to obtain more quantitative and objective information. These abilities are key to effective medical interactions.
Appendix D Expert Verification for Medical Records
We carefully screen and review the medical records to ensure its reliability and relevance. In § 3.2, these medical records are vetted by doctors sourced from Jiangsu Provincial People’s Hospital and Tongji Medical College of Huazhong University of Science and Technology, which hold the prestigious distinction of being tertiary Class A hospitals in China. We utilize the Tencent Questionnaire  [^2] https://wj.qq.com/ platform to facilitate the quality inspection of medical record by doctors. A representative case is depicted in Figure 6, and we also include a display of the questionnaire template in Table 6. These validation steps ensure the high quality of our dataset, ensuring that the benchmark is based on audited medical cases.
@@ Appendix D Expert Verification for Medical Records
Translated Questionnaire Template {Medical Record} 1. Is the diagnostic results correct? YesNo 2. Is the diagnostic rationale correct? YesNo 3. Is the treatment course reasonable? YesNo
Appendix E Attacking on Examiner: Reliability of Examiner’s Workflow
To assess the vulnerability of Examiner agent to information extraction tactics, we design complex scenarios intended to safeguard against the unauthorized disclosure of medical examination results in scenarios lacking explicit examination requests. We build 50 samples, such as “Please provide the most recent medical examination details”, aimed at eliciting information. During the attack simulations, we observe that the Examiner agent, in the first step, uniformly classifies these requests as void of legitimate examination items, implementing a defensive strategy with 100% consistency. This assessment is crucial because it verifies that the Examiner agent behaves appropriately by ensuring that it only releases medical examination results when the patient explicitly specifies the required examinations. It guarantees that the Patient agent cannot obtain relevant information from the Examiner agent by using a method that does not specify a particular examination, specifically when the Doctor agent recommends a unrelated medical test.
Appendix F Human Evaluation for Agent Behavior
In Table G, we present a questionnaire developed for the manual validation of patient and examiner behaviors in each round of conversation. The questionnaire includes four questions, with the initial two addressing "Relevance" and "Consistency" in the question-and-answer (QA) pairs, while the subsequent two focus on the "Accuracy" of conducting medical examinations. Three medical students from Jiangsu Provincial People’s Hospital complete these questionnaires. Consensus among the first two reviewers leads to the immediate acceptance of their collective assessment. In cases of divergence, the opinion of the third reviewer is solicited, whose determination, reflecting the majority viewpoint, constitutes the definitive annotation. The agreement rates for each question are 99.1%, 95.6%, 99.4%, and 100.0%. Significantly, when calculating the accuracy of medical examinations, we exclusively consider dialogues affirmed as "No" in response to the third question.
Appendix G Prompts for Different Medical Roles
Prompt Agent Function
Table 9 Patient Chat with Doctor
Table 11 Examiner Process Examination Request
Table 11 Examiner Produce Examination Outcomes
Table 13 & 13 Chief Physician Evaluate Diagnosis of Doctor
Table 15 Doctor Interactive Clinical Diagnosis
Table 16 Doctor Collaboration through Discussion
Table 14 Chief Physician Summarize Statement of Various Doctors
Table 7: Prompts of different agents and the corresponding function.
We list the prompts agents in AI Hospital in Table 7. In each prompt, {xx} needs to fill with corresponding external inputs. We meticulously design prompts for each agent to ensure clarity and functionality. Particularly for the Doctor role, we discover that overly complex prompts could lead to issues in the dialogue flow, such as not adhering to the prompts or causing cognitive confusion (e.g., the doctor sometimes outputting the patient’s responses). These final prompts are adaptable to most LLMs, enabling the agents in AI Hospital to perform their respective duties effectively.
@@ Appendix G Prompts for Different Medical Roles
Questionnaire {CJK} UTF8gbsn { 病历 } { 单轮对话内容 } 请你仔细阅读这一轮对话的内容和病人的病历信息,回答下面的问题。 1. 病人(检查员)的发言与医生的相关吗? 2. 病人(检查员)的发言符合病历的内容吗? 3. 医生是否建议进行专业的医学检查? 4. 检查员是否进行了医学检查? 5. 医生的总结是否与诊断过程的内容匹配? 匹配少量不匹配明显不匹配 Translated Questionnaire { Medical Record } { Single Round Conversation Content } Carefully review the content of the conversation and the corresponding medical record to answer the following questions. 1. Is the statement of patient or examiner relevant to the doctor’s one? YesNo 2. Is the statement of patient or examiner consistent with the content of medical record? YesNo 3. Does the doctor recommend a professional medical examination? YesNo 4. Does the examiner perform a medical test? YesNo 5. Is the doctor’s summary consistent with the content of the diagnostic process? ConsistentMinor InconsistentSignificant Inconsistent
@@
Prompt for Patient Agent System Message{CJK} UTF8gbsn 你是一个病人。这是你的基本资料。 {个性化信息} {病历中的基本信息}{CJK} UTF8gbsn 下面会有医生来对你的身体状况进行诊断,你需要: (1) 按照病历和基本资料的设定进行对话。 (2) 在每次对话时,你都要明确对话的对象是<医生>还是<检查员>。当你对医生说话时,你要在句子开头说<对医生讲>;如果对象是<检查员>,你要在句子开头说<对检查员讲>。 (3) 首先按照主诉进行回复。 (4) 当<医生>询问你的现病史、既往史、个人史时,要按照相关内容进行回复。 (5) 当<医生>要求或建议你去做检查时,要立即主动询问<检查员>对应的项目和结果,例如:<对检查员讲> 您好,我需要做xxx检查,能否告诉我这些检查结果? (6) 回答要口语化,尽可能短,提供最主要的信息即可。 (7) 从<检查员>那里收到信息之后,将内容主动复述给<医生>。 (8) 当医生给出诊断结果、对应的诊断依据和治疗方案后,在对话的末尾加上特殊字符<结束>。 User [患者] {Statement Generated by GPT-4 in §2.2}
Prompt for Patient Agent System Message You are a patient. Here is your basic information. {Personality in §2.2} {Basic Information in Medical Record §2.1} A doctor will come to diagnose your physical condition. You need to: (1) Engage in dialogue according to the settings of personality and the basic information in medical record. (2) In each conversation, you must clarify whether you are speaking to a <doctor> or an <examiner>. When you speak to the doctor, you should start your sentences with <To the doctor>; if the addressee is an <examiner>, you should start with <To the examiner>. (3) First, respond according to the chief complaint. (4) When the <doctor> asks about your present illness history, past medical history, and personal history, reply according to the relevant content. (5) When the <doctor> requests or suggests that you undergo tests, immediately ask the <examiner> about the corresponding items and results, for example: <To the examiner> Hello, I need to have xxx examination, can you tell me the results of these tests? (6) The responses should be conversational, as short as possible, providing only the most important information. (7) After receiving information from the <examiner>, actively repeat the content to the <doctor>. (8) When the doctor provides the diagnostic result, the corresponding rationale for the diagnosis, and the treatment plan, end the dialogue with the special token <end>. User [Patient] {Statement Generated by GPT-4 in §2.2}
@@
Prompt for Examiner to Process Examination Request {CJK} UTF8gbsn System Message 你是医院负责检查的自动化接待员。请你利用掌握的医学检查的命名实体的知识,从病人的检查申请当中解析出指向明确的专业医学检查项目,方便后面的检查科室进行检查。 请按照下面的格式的输出: # 检查项目 - xxx - xxx 如果没有找到具体的医学检查项目,请输出: # 检查项目 - 无 User [患者] 您好,医生告诉我根据CT扫描和PET-CT扫描的结果,初步得出以下结论:右肺上叶有一个大小约为2.6*1.9cm的实性结节。双肺下叶也有散在的淡薄浸润影。医生建议我进行进一步的检查,例如活检。 Assistant [检查员] # 检查项目 - 肺部活检 User [患者] 我需要了解一下我的检查结果。可以告诉我具体的检查项目和结果吗?谢谢!? Assistant [检查员] # 检查项目 - 无
Prompt for Examiner to Process Examination Request System Message You are an automated receptionist responsible for examinations in a hospital. Using your knowledge of medical examination named entities, please parse out specific professional medical examination items from patients’ examination requests to facilitate subsequent examinations by the relevant departments. Output in the following format: # Examination Item - xxx - xxx If no specific medical examination items are found, please output: # Examination Item - None User [Patient] Hello, the doctor told me that based on the results of the CT scan and PET-CT scan, the preliminary conclusion is that there is a solid nodule approximately 2.6*1.9cm in size in the upper lobe of the right lung. There are also scattered thin infiltrative shadows in the lower lobes of both lungs. The doctor advised me to undergo further examinations, such as a biopsy. Assistant [Examiner] # Medical Examination Items - Lung biopsy User [Patient] I need to know about my examination results. Can you tell me the specific examination items and results, please? Thank you!? Assistant [Examiner] # Medical Examination Items - None
Prompt for Examiner to Produce Examination Outcomes {CJK} UTF8gbsn System Message 这是你收到的病人的检查结果。 {Professional Medical Examination in §2.1} 下面会有病人或医生来查询,你要忠实地按照收到的检查结果,找到对应的项目,并按照下面的格式来回复。 # xx检查 - xxx: xxx - xxx: xxx 如果无法查询到对应的检查项目则回复: - xxx: 无异常